# Provably Robust DPO: Aligning Language Models with Noisy Feedback

### 01 Mar 2024 v1

@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has proven highly effective in aligning Large Language Models (LLMs) with human preferences (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022). In the RLHF pipeline Kaufmann et al. (2023), an LLM is first pre-trained using supervised fine tuning to obtain a reference or SFT policy. A reward model is fit to a dataset of human preferences in the form of a classifier between preferred and rejected responses. Next, an LLM policy is trained using RL algorithms such as proximal policy optimization (PPO) to generate high-reward responses while minimizing a certain notion of divergence from the SFT policy.
@@ 3.1 An Unbiased Loss Function
The BCE loss (9) and the conservative loss (11) have a common drawback – both introduce preference bias in the DPO loss (7). This is due to the following fact:
Fact 3.1 (Logits).
For any $\theta,\theta_{0}\in\mathbb{R}^{d}$ and $\varepsilon\in(0,1/2)$,
[…]
whenever $\mathbb{P}_{\theta}[a\succ a^{\prime}|s]\geq\mathbb{P}_{\theta}[a^{\prime}\succ a|s]$.
That is, whenever $a$ is more likely to be preferred over $a^{\prime}$ without noise, the log-odds of preferring $a$ over $a^{\prime}$ under noisy preferences is at most $(1-\varepsilon)$-th fraction of the corresponding log-odds of that without noise. Ideally, we want the logits to be same for both with and without noise. To this end, we define (un-normalized) preference probabilities
^$\widehat{\mathbb{P}}_{\theta,\varepsilon}[a\succ a^{\prime}|s]=\frac{\sigma(\beta h_{\theta}(s,a,a^{\prime}))^{(1-\varepsilon)}}{\sigma(\beta h_{\theta}(s,a^{\prime},a))^{\varepsilon}}~{},$
which have the same logits as those without noise (see (6)):
$\text{logit}(\widehat{\mathbb{P}}_{\theta,\varepsilon}[a\!\succ\!a^{\prime}|s])\!=\!\frac{\sigma(\!\beta h_{\theta}(s,a,a^{\prime})\!)}{\sigma(\!\beta h_{\theta}(s,a^{\prime},a)\!)}\!=\!\text{logit}(\mathbb{P}_{\theta}[a\!\succ\!a^{\prime}|s]).$
This help us design an unbiased estimator of the loss (7) from noisy preferences as depicted in the following lemma.
@@
For any $\theta,\theta_{0}\!\in\!\mathbb{R}^{d}$, $\varepsilon\!\in\!(0,1/2)$ and noisy preference data $(s,\widetilde{a}_{w},\widetilde{a}_{l})$, define the loss
It holds that $\mathbb{E}_{\varepsilon}\Big{[}\widehat{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})|a_{w},a_{l}\Big{]}=\mathcal{L}(\theta;s,a_{w},a_{l})$.
@@ 6 Experiments
In this section, we provide details about baselines, datasets, and evaluation results. We empirically evaluate $\mathrm{rDPO}$ on two open-ended generation tasks similar to Rafailov et al. (2023): (i) Controlled Sentiment Generation and (ii) Single-turn Dialogue. We compare $\mathrm{rDPO}$ with vanilla $\mathrm{DPO}$ and $\mathrm{cDPO}$ in both tasks. We also compare $\mathrm{rPPO}$ with vanilla $\mathrm{PPO}$ (RLHF) and $\mathrm{cPPO}$ in the sentiment generation task.
@@
Steps $\mathrm{DPO}$ (On clean data) $\mathrm{DPO}$ $\mathrm{cDPO}$ $\mathrm{rDPO}$
200 0.99 $\pm$ 0.03 0.93 $\pm$ 0.26 0.84 $\pm$ 0.36 0.99 $\pm$ 0.00
400 0.99 $\pm$ 0.02 0.72 $\pm$ 0.43 0.82 $\pm$ 0.37 0.99 $\pm$ 0.00
600 0.99 $\pm$ 0.00 0.88 $\pm$ 0.32 0.82 $\pm$ 0.38 0.99 $\pm$ 0.00
800 0.99 $\pm$ 0.00 0.88 $\pm$ 0.32 0.83 $\pm$ 0.36 0.99 $\pm$ 0.00
1000 0.99 $\pm$ 0.02 0.88 $\pm$ 0.32 0.83 $\pm$ 0.37 0.99 $\pm$ 0.00
We then introduce noise into this dataset by randomly flipping preferences with a probability of $\varepsilon=0.4$. For all methods, gpt2-large is employed as the initial policy. For methods in the $\mathrm{DPO}$ family (vanilla $\mathrm{DPO}$, $\mathrm{rDPO}$, $\mathrm{cDPO}$), we optimized the policy for 1000 steps with batch size 16. For methods in the $\mathrm{PPO}$ family (vanilla $\mathrm{PPO}$, $\mathrm{rPPO}$, $\mathrm{cPPO}$), we trained a reward model on preference data for 1000 steps with batch size $16$ and performed policy optimization for 1 epoch over the entire train dataset.
@@
For evaluation, we generate reviews using the final policy and computed rewards using the ground-truth reward model $r^{*}$. The results are presented in Table 1 for the $\mathrm{DPO}$ family and in Table 2 for the $\mathrm{PPO}$ family. For reference, we also train $\mathrm{DPO}$ and $\mathrm{PPO}$ on clean data without any noise. We observe that the performance of $\mathrm{DPO}$ degrades with the introduction of high noise ($\varepsilon=0.4$) in data. However, $\mathrm{rDPO}$ maintains performance across steps, which indicates its robustness to noise. We also observe that $\mathrm{cDPO}$ is not able to mitigate the effect of noise confirming the conclusions of Lemma 3.3. Similar observations are noticed for the $\mathrm{PPO}$ family. In Figure 1, we evaluate average rewards fetched by generations at different sampling temperatures. It is observed that $\mathrm{rDPO}$ and $\mathrm{rPPO}$ achieve the best reward by a significant margin compared to peers in their families.
Table 3: Win-rate on the Anthropic-HH dataset at temparature 1.0.
Method Win rate vs chosen
$\mathrm{sft}$ 0.518
$\mathrm{DPO}$ 0.633
$\mathrm{cDPO}$ ($\varepsilon$ = 0.1) 0.613
$\mathrm{rDPO}$ ($\varepsilon$ = 0.1) 0.644
Figure 2: (a) Evaluation accuracy across steps for different values of label smoothing for $\mathrm{rDPO}$. (b) Evaluation accuracy of $\mathrm{rDPO}$ compared with $\mathrm{DPO}$ and $\mathrm{cDPO}$.
@@
Single-turn Dialogue. In this experiment, each prompt $s$ is a human query and each action $a$ is a helpful response to $s$. We use the Anthropic helpful and harmless dataset (Bai et al., 2022a) as the preference data. We use a supervised fine-tuned gpt2-large model trained on a subset of the chosen preference data as the initial (SFT) policy. We first perform policy optimization using $\mathrm{rDPO}$. As the true noise level in the dataset is unknown, we experiment with different values of $\varepsilon\in\{0.1,0.2,0.3,0.4\}$. We plot the evaluation accuracy of the policy on a subset of the test set across different training steps. This is given by $\frac{1}{m}\sum_{i\in\mathcal{D}_{\text{test}}}\mathds{1}(\widehat{r}_{\theta}(s_{i},a_{w,i})>\widehat{r}_{\theta}(s_{i},a_{l,i}))$, where $\widehat{r}_{\theta}$ is the implicit reward defined by policy $\pi_{\theta}$. The best results are observed with $\varepsilon=0.1$; see Figure 2a. Hence, subsequently, we train $\mathrm{DPO}$ and $\mathrm{cDPO}$ (with label-smoothing $\varepsilon=0.1$) on the same data. We compare the evaluation accuracy of the $\mathrm{DPO}$ family across training steps in Figure 2b. As expected, we observe that $\mathrm{rDPO}$ performs the best.
@@
In this experiment, as we do not have access to any latent reward model, we employ meta-llama/Llama-2-13b-chat-hf 333huggingface.co/meta-llama/Llama-2-13b-chat-hf to compute the win rate of policy generations against the chosen preferences on a representative subset of the test dataset. The win-rates can be seen in Table 3. We observe a slightly better performance of $\mathrm{rDPO}$ compared to $\mathrm{DPO}$. We conjecture that this is due to low noise present in train set as evidenced by the best results of $\mathrm{rDPO}$ at $\varepsilon=0.1$.
@@ Appendix A Missing Details
A.1 Proof of Fact 3.1
This follows from Appendix C.2.1 of Chowdhury et al. (2023) by substituting their probability of no-flip $\sigma(\varepsilon)$ by our probability of no-flip $1-\varepsilon$.
A.2 Proof of Lemma 3.2
@@
Denote $z_{i}=(s_{i},\widetilde{a}_{w,i},\widetilde{a}_{l,i})$. Then, our de-biased loss function (12) can be re-written as
@@
$\widehat{\Sigma}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\phi(s_{i},a_{w,i})-\phi(s_{i},a_{w,i})\right)\left(\phi(s_{i},a_{w,i})-\phi(s_{i},a_{w,i})\right)^{\top},$
@@
Since $a,a^{\prime}$ are indpendent samples from policy $\pi(\cdot|s)$, it holds that

### 12 Apr 2024 v2

@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has proven highly effective in aligning Large Language Models (LLMs) with human preferences (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022). In the RLHF pipeline(Kaufmann et al., 2023), an LLM is first pre-trained using supervised fine tuning to obtain a reference or SFT policy. A reward model is fit to a dataset of human preferences in the form of a classifier between preferred and rejected responses. Next, an LLM policy is trained using RL algorithms such as proximal policy optimization (PPO) to generate high-reward responses while minimizing a certain notion of divergence from the SFT policy.
@@ 3.1 An Unbiased Loss Function
The BCE loss (9) and the conservative loss (11) have a common drawback – both introduce bias in the DPO loss (7). This is due to the fact that
$\mathbb{E}\left[\ell(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\right]\neq\mathcal{L}(\theta;s,a_{w},a_{l}),\,\ell\in\{\mathcal{L}_{\varepsilon},\bar{\mathcal{L}}_{\varepsilon}\}~{}.$
It also holds that (Chowdhury et al., 2023)
$\text{logit}(\mathbb{P}_{\theta,\varepsilon}[a\succ a^{\prime}|s])\neq\text{logit}(\mathbb{P}_{\theta}[a\succ a^{\prime}|s]).$
That is, the log-odds of preferring $a$ over $a^{\prime}$ under noisy preferences is different from that without noise, which introduces a bias in preferences. Ideally, we want the logits to be same for both with and without noise. To this end, we define (un-normalized) preference probabilities
^$\widehat{\mathbb{P}}_{\theta,\varepsilon}[a\succ a^{\prime}|s]=\frac{\sigma(\beta h_{\theta}(s,a,a^{\prime}))^{(1-\varepsilon)}}{\sigma(\beta h_{\theta}(s,a^{\prime},a))^{\varepsilon}}~{}.$
these have the same logits as those without noise, since
$\text{logit}(\widehat{\mathbb{P}}_{\theta,\varepsilon}[a\!\succ\!a^{\prime}|s])$ $\!=\!\log\left(\frac{\sigma(\!\beta h_{\theta}(s,a,a^{\prime})\!)}{\sigma(\!\beta h_{\theta}(s,a^{\prime},a)\!)}\right)$
$\!=\!\text{logit}(\mathbb{P}_{\theta}[a\!\succ\!a^{\prime}|s]).$
This motivates us to define the loss function:
This loss is an unbiased estimator of the DPO loss (7) under noisy preferences as stated in the following lemma.
@@
For any $\theta,\theta_{0}\!\in\!\mathbb{R}^{d}$, $\varepsilon\!\in\!(0,1/2)$, we have
$\mathbb{E}_{\varepsilon}\Big{[}\widehat{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})|a_{w},a_{l}\Big{]}=\mathcal{L}(\theta;s,a_{w},a_{l})~{}.$
@@
Variance of rDPO loss. Along with unbiasedness, it is also desirable to have bounded variance of the estimator. To this end, consider the un-normalized rDPO loss $(1-2\varepsilon)\widehat{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})$, which yields the same loss-minimizing policy as in (13). It has a variance $\varepsilon(1-\varepsilon)\left[\mathcal{L}(\theta;s,a_{w},a_{l})-\mathcal{L}(\theta;s,a_{l},a_{w})\right]^{2}$. For Neural policy class of the form (4) and for bounded $f_{\theta}$, the variance is bounded by $C\varepsilon(1-\varepsilon)$ for some constant $C>0$. Since $\varepsilon\leq 1/2$, the variance is bounded by $C/4$.
@@ 4.2 Performance Bounds of Learned Policy
Data efficiency of rDPO under a given noise level. We can obtain a bound on sample complexity of $\mathrm{rDPO}$ for a given noise level and permissible sub-optimality gap. For instance, if $\widehat{\Sigma}$ is invertible, then training rDPO on $n\geq\frac{\kappa d}{\Delta^{2}\gamma^{2}\beta^{2}(1-2\varepsilon)^{2}}$ samples, we can ensure a sub-optimality gap $\leq\Delta$ for the aligned model. In contrast, when samples are clean ($\varepsilon=0$), then training vanilla $\mathrm{DPO}$ on $n\geq\frac{\kappa d}{\Delta^{2}\gamma^{2}\beta^{2}}$ samples, we can ensure a suboptimality gap $\leq\Delta$. Thus, under the presence of noise, $\mathrm{rDPO}$ needs roughly $\frac{1}{(1-2\varepsilon)^{2}}$ times the samples that $\mathrm{DPO}$ needs under clean data. The higher the noise level, the higher the number of samples needed for $\mathrm{rDPO}$.
@@ 6 Experiments
In this section, we provide details about baselines, datasets, and evaluation results. We empirically evaluate $\mathrm{rDPO}$ on two open-ended generation tasks similar to Rafailov et al. (2023): (i) Controlled Sentiment Generation and (ii) Single-turn Dialogue. We compare $\mathrm{rDPO}$ with vanilla $\mathrm{DPO}$ and $\mathrm{cDPO}$ in both tasks. In the sentiment generation task, we also include $\mathrm{SLiC}$ (Zhao et al., 2023) and $\mathrm{IPO}$ (Azar et al., 2023) as baselines. Furthermore, we compare $\mathrm{rPPO}$ with vanilla $\mathrm{PPO}$ (RLHF) and $\mathrm{cPPO}$ in this task.
@@
Steps $\mathrm{DPO}$ (On clean data) $\mathrm{DPO}$ $\mathrm{cDPO}$ $\mathrm{IPO}$ $\mathrm{SLiC}$ $\mathrm{rDPO}$
200 0.99 $\pm$ 0.03 0.93 $\pm$ 0.26 0.84 $\pm$ 0.36 0.85 $\pm$ 0.35 0.94 $\pm$ 0.22 0.99 $\pm$ 0.00
400 0.99 $\pm$ 0.02 0.72 $\pm$ 0.43 0.82 $\pm$ 0.37 0.83 $\pm$ 0.37 0.88 $\pm$ 0.31 0.99 $\pm$ 0.00
600 0.99 $\pm$ 0.00 0.88 $\pm$ 0.32 0.82 $\pm$ 0.38 0.84 $\pm$ 0.36 0.90 $\pm$ 0.29 0.99 $\pm$ 0.00
800 0.99 $\pm$ 0.00 0.88 $\pm$ 0.32 0.83 $\pm$ 0.36 0.83 $\pm$ 0.37 0.89 $\pm$ 0.30 0.99 $\pm$ 0.00
1000 0.99 $\pm$ 0.02 0.88 $\pm$ 0.32 0.83 $\pm$ 0.37 0.82 $\pm$ 0.38 0.90 $\pm$ 0.29 0.99 $\pm$ 0.00
We then introduce noise into this dataset by randomly flipping preferences with a probability of $\varepsilon=0.4$. For all methods, gpt2-large is employed as the initial policy. For methods in the $\mathrm{DPO}$ family (vanilla $\mathrm{DPO}$, $\mathrm{rDPO}$, $\mathrm{cDPO}$), we optimized the policy for 1000 steps with batch size 16. We do the same for $\mathrm{IPO}$ and $\mathrm{SLiC}$. For methods in the $\mathrm{PPO}$ family (vanilla $\mathrm{PPO}$, $\mathrm{rPPO}$, $\mathrm{cPPO}$), we trained a reward model on preference data for 1000 steps with batch size $16$ and performed policy optimization for 1 epoch over the entire train dataset.
@@
For evaluation, we generate reviews using the final policy and computed rewards using the ground-truth reward model $r^{*}$. The results are presented in Table 1 for the $\mathrm{DPO}$ family and in Table 2 for the $\mathrm{PPO}$ family. For reference, we also train $\mathrm{DPO}$ and $\mathrm{PPO}$ on clean data without any noise. We observe that the performance of $\mathrm{DPO}$ degrades with the introduction of high noise ($\varepsilon=0.4$) in data. $\mathrm{IPO}$ and $\mathrm{SLiC}$ also suffers significantly due to noisy preferences. However, $\mathrm{rDPO}$ maintains performance across steps, which indicates its robustness to noise. We also observe that $\mathrm{cDPO}$ is not able to mitigate the effect of noise confirming the conclusions of Lemma 3.2. Similar observations are noticed for the $\mathrm{PPO}$ family. In Figure 1, we evaluate average rewards fetched by generations at different sampling temperatures. It is observed that $\mathrm{rDPO}$ and $\mathrm{rPPO}$ achieve the best reward by a significant margin compared to peers in their families.
Table 3: Percentage Improvement on win-rate vs chosen response over the initial SFT policy
Method Improvement over SFT (%)
gpt2-large Llama-2-7b
$\mathrm{DPO}$ 22.20 45.78
$\mathrm{cDPO}$ ($\varepsilon$ = 0.1) 18.34 39.16
$\mathrm{rDPO}$ ($\varepsilon$ = 0.1) 24.32 51.20
@@
Single-turn Dialogue. In this experiment, each prompt $s$ is a human query and each action $a$ is a helpful response to $s$. We use the Anthropic helpful and harmless dataset (Bai et al., 2022a) as the preference data. We use a supervised fine-tuned gpt2-large model trained on a subset of the chosen preference data as the initial (SFT) policy. We first perform policy optimization using $\mathrm{rDPO}$. As the true noise level in the dataset is unknown, we experiment with different values of $\varepsilon\in\{0.1,0.2,0.3,0.4\}$. We plot the evaluation accuracy of the policy on a subset of the test set across different training steps. This is given by $\frac{1}{m}\sum_{i\in\mathcal{D}_{\text{test}}}\mathds{1}(\widehat{r}_{\theta}(s_{i},a_{w,i})>\widehat{r}_{\theta}(s_{i},a_{l,i}))$, where $\widehat{r}_{\theta}$ is the implicit reward defined by policy $\pi_{\theta}$. We observed the best results with $\varepsilon=0.1$. Subsequently, we train $\mathrm{DPO}$ and $\mathrm{cDPO}$ (with label-smoothing $\varepsilon=0.1$) on the same data.
@@
In this experiment, as we do not have access to any latent reward model, we employ meta-llama/Llama-2-13b-chat-hf 333huggingface.co/meta-llama/Llama-2-13b-chat-hf to compute the win rate of policy generations against the chosen preferences on a representative subset of the test dataset. Next, to demonstrate that of our method generalizes to bigger models, we repeat this experiment with Llama-2-7b as the policy model and GPT-4 as the evaluation model. The win-rates for both experiments are tabulated in Table 3. In both cases, we observe that $\mathrm{rDPO}$ performs significantly better than $\mathrm{DPO}$ and $\mathrm{cDPO}$.
@@ Appendix A Missing Details
A.1 Proof of Lemma 3.1
@@
A.2 Variance of rDPO loss
First, define the un-normalized rDPO loss
~$\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l}):=(1-2\varepsilon)\widehat{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})=(1-\varepsilon)\mathcal{L}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})-\varepsilon\mathcal{L}(\theta;s,\widetilde{a}_{l},\widetilde{a}_{w})~{}.$
Its variance is given by
$\text{Var}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\right]$ $=\mathbb{E}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})^{2}\right]-\mathbb{E}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\right]^{2}~{}.$
From Lemma 3.1, we have
$\mathbb{E}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\right]=(1-2\varepsilon)\mathcal{L}(\theta;s,a_{w},a_{l})~{}.$
Furthermore, we have
$\mathbb{E}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})^{2}\right]$ $=(1-\varepsilon)^{2}\mathbb{E}\left[\mathcal{L}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})^{2}\right]+\varepsilon^{2}\mathbb{E}\left[\mathcal{L}(\theta;s,\widetilde{a}_{l},\widetilde{a}_{w})^{2}\right]-2\varepsilon(1-\varepsilon)\mathbb{E}\left[\mathcal{L}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\mathcal{L}(\theta;s,\widetilde{a}_{l},\widetilde{a}_{w})\right]~{}.$
Now observe that
$\mathbb{E}\left[\mathcal{L}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})^{2}\right]=(1-\varepsilon)\mathcal{L}(\theta;s,a_{w},a_{l})^{2}+\varepsilon\mathcal{L}(\theta;s,a_{l},a_{w})^{2}~{},$
$\mathbb{E}\left[\mathcal{L}(\theta;s,\widetilde{a}_{l},\widetilde{a}_{w})^{2}\right]=(1-\varepsilon)\mathcal{L}(\theta;s,a_{l},a_{w})^{2}+\varepsilon\mathcal{L}(\theta;s,a_{w},a_{l})^{2}~{},$
$\mathbb{E}\left[\mathcal{L}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\mathcal{L}(\theta;s,\widetilde{a}_{l},\widetilde{a}_{w})\right]=\mathcal{L}(\theta;s,a_{w},a_{l})\mathcal{L}(\theta;s,a_{l},a_{w})~{}.$
Combining all these, we get
$\mathbb{E}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})^{2}\right]=(1-3\varepsilon+3\varepsilon^{2})\mathcal{L}(\theta;s,a_{w},a_{l})^{2}+\varepsilon(1-\varepsilon)\mathcal{L}(\theta;s,a_{l},a_{w})^{2}-2\varepsilon(1-\varepsilon)\mathcal{L}(\theta;s,a_{w},a_{l})\mathcal{L}(\theta;s,a_{l},a_{w})~{}.$
Therefore, the variance of the un-normalized rDPO loss is given by
$\text{Var}\left[\widetilde{\mathcal{L}}_{\varepsilon}(\theta;s,\widetilde{a}_{w},\widetilde{a}_{l})\right]$
$=\left[(1-3\varepsilon+3\varepsilon^{2})-(1-2\varepsilon)^{2}\right]\mathcal{L}(\theta;s,a_{w},a_{l})^{2}+\varepsilon(1-\varepsilon)\mathcal{L}(\theta;s,a_{l},a_{w})^{2}-2\varepsilon(1-\varepsilon)\mathcal{L}(\theta;s,a_{w},a_{l})\mathcal{L}(\theta;s,a_{l},a_{w})$
$=\varepsilon(1-\varepsilon)\left[\mathcal{L}(\theta;s,a_{w},a_{l})^{2}+\mathcal{L}(\theta;s,a_{l},a_{w})^{2}-2\mathcal{L}(\theta;s,a_{w},a_{l})\mathcal{L}(\theta;s,a_{l},a_{w})\right]$
$=\varepsilon(1-\varepsilon)\left[\mathcal{L}(\theta;s,a_{w},a_{l})-\mathcal{L}(\theta;s,a_{l},a_{w})\right]^{2}.$
@@ A.2 Variance of rDPO loss
Denote $z_{i}=(s_{i},\widetilde{a}_{w,i},\widetilde{a}_{l,i})$. Then, our de-biased loss function (12) can be re-written as [^4] We ignore the normalization by $1-2\varepsilon$, since it doesn’t affect the minimizer of the loss.
@@
$\widehat{\Sigma}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\phi(s_{i},a_{w,i})-\phi(s_{i},a_{l,i})\right)\left(\phi(s_{i},a_{w,i})-\phi(s_{i},a_{l,i})\right)^{\top},$
@@
Since $a,a^{\prime}$ are independent samples from policy $\pi(\cdot|s)$, it holds that